# Gröbner Basis Cryptanalysis of Ciminion and Hydra

### v1

@@ 2.5 The Eigenvalue Method
$\mathcal{O}\left(n\cdot d^{\omega}\right),$ (27)
where $d=\dim_{K}\left(P/I\right)$. In addition, if we are only interested in solutions over a finite field $\mathbb{F}_{q}$, then factorization of the characteristic polynomials can be done via the GCD with the field equation whose complexity was given in Equation 1.
Note that the probabilistic FGLM algorithm [FGHR14, Proposition 3] has polylogarithmic complexity $\tilde{\mathcal{O}}\left(d^{\omega}\right)$ for the construction of a LEX Gröbner basis of $I$ if $I$ is in shape position, and polylogarithmic complexity $\tilde{\mathcal{O}}\left(n\cdot d^{\omega}\right)$ for the construction of a LEX Gröbner basis of $\sqrt{I}$, which applies to any ideal not in shape position. This is no coincidence, the probabilistic FGLM algorithm constructs a LEX Gröbner basis via the multiplication matrix for the last variable. Hence, for both methods the complexity is dominated by $\mathcal{O}\left(d^{\omega}\right)$.
@@ 3.1 A DRL Gröbner Basis for Ciminion
For simplicity let us denote the matrix in the $i$th Ciminion round by $\mathbf{A}_{i}\in\mathbb{F}_{q}^{3\times 3}$, then
$\mathbf{A}_{i}^{-1}=\begin{pmatrix}0&1&-c_{4}^{(i)}\\ -1&0&1\\ 1&0&0\end{pmatrix},$ (28)
where $c_{4}^{(i)}\notin\mathbb{F}_{q}\setminus\{0,1\}$.
@@
Let $\mathbb{F}_{q}$ be a finite field, and let
$\mathcal{F}_{\texttt{Ciminion}}=\{\mathbf{f}^{(1)},\dots,\mathbf{f}^{(r_{C}+r_{E})}\}\subset\mathbb{F}_{q}\Big{[}y_{1},y_{2},\mathbf{x}_{C}^{(1)},\allowbreak\dots,\mathbf{x}_{C}^{(r_{C}-1)},\allowbreak\mathbf{x}_{E}^{(1)},\dots,\mathbf{x}_{E}^{(r_{E})},x\Big{]}$
be a Ciminion polynomial system. For the DRL term order $y_{1}>y_{2}>\mathbf{x}_{C}^{(1)}>\ldots>\mathbf{x}_{C}^{(r_{C}-1)}>\mathbf{x}_{E}^{(1)}>\mathbf{x}_{E}^{(r_{E})}>x$, the following procedure computes a DRL Gröbner basis for Ciminion:
@@
    For every round $1\leq i\leq r_{C}+r_{E}$ compute $\mathbf{g}^{(i)}=\mathbf{A}_{i}^{-1}\mathbf{g}_{i}$.
@@
$\mathbf{g}^{(1)}$ =($\ y_{1}\\ y_{2}+\aleph\cdot y_{1}\end{pmatrix}+\mathbf{A}_{1}^{-1}\left(\mathbf{c}^{(1)}-\mathbf{x}_{C}^{(1)}\right),$ (29)
$\mathbf{g}^{(i)}$ $=\begin{pmatrix}x_{C,1}^{(i-1)}\\[5.0pt] x_{C,2}^{(i-1)}\\[5.0pt] x_{C,3}^{(i-1)}+x_{C,1}^{(i-1)}\cdot x_{C,2}^{(i-1)}\end{pmatrix}+\mathbf{A}_{i}^{-1}\left(\mathbf{c}^{(i)}-\begin{pmatrix}x_{C,1}^{(i)}\\ x_{C,2}^{(i)}\\ x_{C,3}^{(i)}\end{pmatrix}\right),$ (30)
$\mathbf{g}^{(r_{C}+r_{E})}$ $=\begin{pmatrix}x_{E,1}^{(r_{E})}\\[5.0pt] x_{E,2}^{(r_{E})}\\[5.0pt] x_{E,3}^{(r_{E})}+x_{E,1}^{(r_{E})}\cdot x_{C,2}^{(r_{E})}\end{pmatrix}+\mathbf{A}_{i}^{-1}\left(\mathbf{c}^{(r_{E})}-\begin{pmatrix}c_{1}-p_{1}\\ c_{2}-p_{2}\\ x\end{pmatrix}\right),$ (31)
and analog for the equations involving the $\mathbf{x}_{E}^{(i)}$’s. Hence, $\mathcal{L}$ contains $\mathbf{g}^{(1)}$ and the first two components of all remaining $\mathbf{g}^{(i)}$’s.
For Gaussian elimination on $\mathcal{L}$ let us start in the last round $i=r_{C}+r_{E}$. By construction and Equation 28 the linear polynomials with leading monomials $x_{E,1}^{(r_{E})}$, $x_{E,2}^{(r_{E})}$ contain a linear term in $x$ and a constant term. Moving to $i=r_{C}+r_{E}-1$ next, then the linear polynomials from $\mathbf{g}^{(r_{C}+r_{E}-1)}$ depend on $\mathbf{x}_{E}^{(r_{C}+r_{E}-1)}$ and $\mathbf{x}_{E}^{(r_{C}+r_{E})}$, via Gaussian elimination the variables $x_{E,1}^{(r_{E})}$ and $x_{E,2}^{(r_{E})}$ get eliminated. Thus, the linear polynomials with leading monomials in $x_{E,1}^{(r_{E}-1)}$, $x_{E,2}^{(r_{E}-1)}$ can only depend on $x_{E,3}^{(r_{E})}$ and $x$. Moreover, by Equation 28 the coefficients for $x_{E,3}^{(r_{E})}$ are non-zero. Via a downwards induction we can then conclude that the linear polynomials with leading monomials $x_{\ast,1}^{(i)}$, $x_{\ast,2}^{(i)}$ only depend on the variables $x_{\ast,3}^{(i+1)},\dots,x_{E,3}^{(r_{E})},x$ and that the coefficients of $x_{\ast,3}^{(i+1)}$ are non-zero. The first round needs special care since the first row of $\mathbf{A}_{1}^{-1}$ is non-zero on the variable $x_{C,2}^{(1)}$. (If it were zero on $x_{C,1}^{(1)}$ and $x_{C,2}^{(1)}$, the first polynomial in $\mathbf{g}^{(1)}$ would have leading monomial $x_{C,3}^{(1)}$ and the proof would be finished.) Expansion of Equation 29 yields
$\mathbf{g}^{(1)}=\begin{pmatrix}\aleph\\ y_{1}\\ y_{2}+\aleph\cdot y_{1}\end{pmatrix}+\mathbf{A}_{1}^{-1}\mathbf{c}_{1}-\begin{pmatrix}x_{C,2}^{(1)}-c_{4}^{(1)}\cdot x_{C,3}^{(1)}\\ -x_{C,1}^{(1)}+x_{C,3}^{(1)}\\ x_{C,1}^{(1)}\end{pmatrix}.$
Producing a polynomial with leading monomial $y_{2}$ is straight-forward, we simply have to take $g_{3}^{(1)}-\aleph\cdot g_{2}^{(1)}$. Note that there is second polynomial with leading monomial $x_{C,2}^{(1)}$ after inversion of the matrices
$g_{2}^{(2)}=x_{C,2}^{(1)}+\left(\mathbf{A}_{2}^{-1}\mathbf{c}^{(2)}\right)_{2}-\left(-x_{C,1}^{(2)}+x_{C,3}^{(2)}\right).$
@@
$g_{1}^{(1)}+g_{2}^{(2)}=\aleph+\left(\mathbf{A}_{1}^{-1}\mathbf{c}^{(1)}\right)_{1}+c_{4}^{(1)}\cdot x_{C,3}^{(1)}+\left(\mathbf{A}_{2}^{-1}\mathbf{c}^{(2)}\right)_{2}+x_{C,1}^{(2)}-x_{C,3}^{(2)},$
and by the choice of our DRL term order and $c_{4}^{(1)}\neq 0$ this polynomial has leading monomial $x_{C,3}^{(1)}$. Obviously, after Gaussian elimination of $\mathcal{L}$ the terms of $g_{1}^{(1)}+g_{2}^{(2)},g_{2}^{(1)},g_{3}^{(1)}-\aleph\cdot g_{2}^{(1)}$ after the leading monomials will also only depend on the variables $x_{C,3}^{(2)},\dots,x_{E,3}^{(r_{E})},x$.
@@
$y_{1}+\mathcal{A}_{1}\left(x_{C,3}^{(2)},\dots,x_{E,3}^{(r_{C}+r_{E}-1)},x\right)$ $=0,$
$y_{2}+\mathcal{A}_{2}\left(x_{C,3}^{(2)},\dots,x_{E,3}^{(r_{C}+r_{E}-1)},x\right)$ $=0,$
$x_{E,3}^{(1)}+\mathcal{A}_{3}\left(x_{C,3}^{(2)},\dots,x_{E,3}^{(r_{C}+r_{E}-1)},x\right)$ $=0,$
$x_{\ast,1}^{(i)}+c_{4}^{(i)}\cdot x_{\ast,3}^{(i+1)}+\ldots$ $=0,$
$x_{\ast,2}^{(i)}-x_{\ast,3}^{(i+1)}+\ldots$ $=0,$
$x_{E,1}^{(r_{E})}+c_{4}^{(r_{C}+r_{E})}\cdot x+\ldots$ $=0,$
$x_{E,2}^{(r_{E})}-x+\ldots$ $=0,$
where $\mathcal{A}_{1}$, $\mathcal{A}_{2}$ and $\mathcal{A}_{3}$ are affine polynomials. So, we can directly conclude from Equation 30 that
$\hat{g}_{3}^{(i)}=\left(g_{3}^{(i)}\mod(\mathcal{L})\right)\in\mathbb{F}_{q}\Big{[}x_{C,3}^{(2)},\dots,x_{E,3}^{(r_{C}+r_{E}-1)},x\Big{]},$
where $2\leq i\leq r_{C}+r_{E}$, has leading monomial ${x_{\ast,3}^{(i)}}^{2}$ and respectively $x^{2}$ for the last round.
@@
Note that if we only keep the $r_{E}+r_{C}-1$ quadratic polynomials of $\mathcal{G}$, then we still have a fully determined polynomial system in $r_{E}+r_{C}-1$ variables as well as a Gröbner basis. Therefore, we call the quadratic polynomials of $\mathcal{G}$ the downsized Ciminion polynomial system.
@@
$\begin{split}f_{1}&=x_{1}-\tilde{f}_{n}(x_{n}),\\ &\dots\\ f_{n-1}&=x_{n-1}-\tilde{f}_{n-1}(x_{n}),\\ f_{n}&=\tilde{f}_{n}(x_{n}).\end{split}$ (32)
Note that the LEX Gröbner basis of a radical ideal can be brought into shape position for some ordering of the variables, see [KR00, Theore 3.7.25].
@@
In the situation of LABEL:Th:_Ciminion_Gr\"obner_basis, let $\mathcal{G}_{\texttt{Ciminion}}$ denote the Ciminion DRL Gröbner basis.
@@
    $\dim_{\mathbb{F}_{q}}\left(\mathcal{F}_{\texttt{Ciminion}}\right)=2^{r_{C}+r_{E}-1}$.
@@
    Let $I\subset\overline{\mathbb{F}_{q}}\Big{[}y_{1},y_{2},\mathbf{x}_{C}^{(1)},\allowbreak\dots,\mathbf{x}_{C}^{(r_{C}-1)},\allowbreak\mathbf{x}_{E}^{(1)},\dots,\mathbf{x}_{E}^{(r_{E})},x\Big{]}[x_{0}]$ be a homogeneous ideal such that $\mathcal{Z}_{+}(I)\neq\emptyset$ and $\mathcal{G}^{\text{hom}}\subset I$. Then $I$ is in generic coordinates.
@@
    $\operatorname{sd}_{DRL}\left(\mathcal{F}_{\texttt{Ciminion}}\right)\leq r_{C}+r_{E}$.
@@
    The LEX Gröbner basis for Ciminion under $y_{1}>y_{2}>\mathbf{x}_{C}^{(1)}>\ldots>\mathbf{x}_{C}^{(r_{C}-1)}>\mathbf{x}_{E}^{(1)}>\mathbf{x}_{E}^{(r_{E})}>x$ is in $x$-shape position.
@@
(1), (2) and (3) are consequences of the DRL Gröbner basis from LABEL:Th:_Ciminion_Gr\"obner_basis and application of Theorems 2.5 and 2.6. For the last claim we take a look at the shape of the DRL Gröbner basis
$g_{3}^{(2)}$ $=\mathcal{Q}_{2}\left(x_{C,3}^{(2)},\dots,x_{E,3}^{(i)},x\right)-\mathcal{A}_{2}\left(x_{C,3}^{(2)},\dots,x_{E,3}^{(r_{E})},x\right),$
$\dots$
$g_{3}^{(i)}$ $=\mathcal{Q}_{i}\left(x_{\ast,3}^{(i)},\dots,x_{E,3}^{(r_{E})},x\right)-\mathcal{A}_{i}\left(x_{\ast,3}^{(i-1)},\dots,x_{E,3}^{(r_{E})},x\right),$
$\dots$
$g_{3}^{(r_{C}+r_{E})}$ $=\mathcal{Q}\left(x\right)-\mathcal{A}_{r_{C}+r_{E}}\left(x_{E,3}^{(r_{E})},x\right),$
where the $\mathcal{Q}_{i}$’s are quadratic polynomials which contain the monomial ${x_{\ast,3}^{(i)}}^{2}$ and respectively $x^{2}$ for the last round, and the $\mathcal{A}_{i}$’s are affine polynomials that contain the monomial $x_{E,3}^{(i-1)}$ for all $i\geq 3$ and without any restrictions on $\mathcal{A}_{2}$. Then by iterative substitution of $g_{3}^{(r_{C}+r_{E})}$ into $g_{3}^{(r_{C}+r_{E}-1)}$, $g_{3}^{(r_{C}+r_{E}-1)}$ into $g_{3}^{(r_{C}+r_{E}-2)}$, etc. we produce a LEX Gröbner basis in $x$-shape position. ∎

### v4

@@ 2.5.1 Solving Structured Polynomial Systems Over Finite Fields
$\mathcal{O}\left(\begin{rcases}\begin{dcases}\sum_{i=1}^{\frac{n}{2}}N^{\frac{n}{2}-i}\cdot\Big{(}2^{\omega\cdot(2\cdot i-1)+1}+\mathcal{D}\big{(}2^{2\cdot i},q\big{)}\Big{)},&2\mid n,\\ \sum_{i=1}^{\frac{n-1}{2}}N^{\frac{n-1}{2}-i}\cdot\Big{(}2^{\omega\cdot 2\cdot i+1}+\mathcal{D}\big{(}2^{2\cdot i+1},q\big{)}\Big{)},&2\nmid n\end{dcases}\end{rcases}\right)$ (24)
$=\mathcal{O}\left(\begin{rcases}\begin{dcases}2^{\omega+1}\cdot\frac{2^{\omega\cdot n}-N^{\frac{n}{2}}}{2^{2\cdot\omega}-N}+\sum_{i=1}^{\frac{n}{2}}N^{\frac{n}{2}-i}\cdot\mathcal{D}\big{(}2^{2\cdot i},q\big{)},&2\mid n,\\ 2^{2\cdot\omega+1}\cdot\frac{2^{\omega\cdot(n-1)}-N^{\frac{n-1}{2}}}{2^{2\cdot\omega}-N}+\sum_{i=1}^{\frac{n-1}{2}}N^{\frac{n-1}{2}-i}\cdot\mathcal{D}\big{(}2^{2\cdot i+1},q\big{)},&2\nmid n\end{dcases}\end{rcases}\right)$ (25)
field operations.
@@ 3.1 A DRL Gröbner Basis for Ciminion
For simplicity let us denote the matrix in the $i$th Ciminion round by $\mathbf{A}_{i}\in\mathbb{F}_{q}^{3\times 3}$, see Equation 2, then
$\mathbf{A}_{i}=\begin{pmatrix}0&0&1\\ 1&c_{4}^{(i)}&c_{4}^{(i)}\\ 0&1&1\end{pmatrix},\qquad\mathbf{A}_{i}^{-1}=\begin{pmatrix}0&1&-c_{4}^{(i)}\\ -1&0&1\\ 1&0&0\end{pmatrix},$ (26)
where $c_{4}^{(i)}\in\mathbb{F}_{q}\setminus\{0,1\}$.
@@
Let $\mathbb{F}_{q}$ be a finite field, and let $\mathcal{F}_{\texttt{Ciminion}}=\left\{\mathbf{f}^{(1)},\dots,\mathbf{f}^{(r)}\right\}\subset\mathbb{F}_{q}\Big{[}y_{1},y_{2},\mathbf{x}^{(1)},\allowbreak\dots,\mathbf{x}^{(r-1)},x\Big{]}$ be a Ciminion polynomial system. For the DRL term order $y_{1}>y_{2}>\mathbf{x}^{(1)}>\ldots>\mathbf{x}^{(r-1)}>x$, the following procedure computes a DRL Gröbner basis for Ciminion:
@@
    For every round $1\leq i\leq r$ compute $\mathbf{g}^{(i)}=\mathbf{A}_{i}^{-1}\mathbf{f}_{i}$.
@@
$\mathbf{g}^{(1)}$ =($\ y_{1}\\ \aleph\cdot y_{1}+y_{2}\end{pmatrix}+\mathbf{A}_{1}^{-1}\left(\mathbf{c}^{(1)}-\mathbf{x}^{(1)}\right),$ (27)
$\mathbf{g}^{(i)}$ $=\begin{pmatrix}x_{1}^{(i-1)}\\ x_{2}^{(i-1)}\\ x_{3}^{(i-1)}+x_{1}^{(i-1)}\cdot x_{2}^{(i-1)}\end{pmatrix}+\mathbf{A}_{i}^{-1}\left(\mathbf{c}^{(i)}-\mathbf{x}^{(i)}\right),$ (28)
$\mathbf{g}^{(r)}$ $=\begin{pmatrix}x_{1}^{(r-1)}\\ x_{2}^{(r-1)}\\ x_{3}^{(r-1)}+x_{1}^{(r-1)}\cdot x_{2}^{(r-1)}\end{pmatrix}+\mathbf{A}_{r}^{-1}\left(\mathbf{c}^{(r)}-\begin{pmatrix}c_{1}-p_{1}\\ c_{2}-p_{2}\\ x\end{pmatrix}\right).$ (29)
Hence, $\mathcal{L}$ contains $\mathbf{g}^{(1)}$ and the first two components of all remaining $\mathbf{g}^{(i)}$’s.
For Gaussian elimination on $\mathcal{L}$ let us start in the last round $i=r$. By Equations 26 and 29 the linear polynomials with leading monomials $x_{1}^{(r-1)}$, $x_{2}^{(r-1)}$ contain a linear term in $x$ and a constant term. Moving to $i=r-1$ next the linear polynomials from $\mathbf{g}^{(r-1)}$ depend on $\mathbf{x}^{(r-2)}$ and $\mathbf{x}^{(r-1)}$, and via Gaussian elimination the variables $x_{1}^{(r-1)}$ and $x_{2}^{(r-1)}$ get eliminated. Thus, the linear polynomials with leading monomials in $x_{1}^{(r-2)}$ and $x_{2}^{(r-2)}$ can only depend on $x_{3}^{(r-1)}$ and $x$. Moreover, by Equation 26 the coefficients for $x_{3}^{(r-1)}$ are non-zero. Via a downwards induction we can then conclude for $12\leq i\leq r-$ 1 that after Gaussian elimination
$\operatorname{LM}_{DRL}\left(g_{1}^{(i)}\right)$ $=x_{1}^{(i-1)},$
$\operatorname{LM}_{DRL}\left(g_{2}^{(i)}\right)$ $=x_{2}^{(i-1)},$
$g_{1}^{(i)},g_{2}^{(i)}$ $\in\mathbb{F}_{q}\big{[}x_{3}^{(i)},\dots,x_{3}^{(r-1)},x\big{]},$
and that the coefficients of $x_{3}^{(i)}$ in $g_{1}^{(i)}$ and $g_{2}^{(i)}$ are non-zero.
The first round needs special care since the first row of $\mathbf{A}_{1}^{-1}$ is non-zero on the variable $x_{2}^{(1)}$. (If it were zero on $x_{1}^{(1)}$ and $x_{2}^{(1)}$, the first polynomial in $\mathbf{g}^{(1)}$ would have leading monomial $x_{3}^{(1)}$, and we could proceed.) Expansion of Equation 27 yields
$\mathbf{g}^{(1)}=\begin{pmatrix}\aleph\\ y_{1}\\ \aleph\cdot y_{1}+y_{2}\end{pmatrix}+\mathbf{A}_{1}^{-1}\mathbf{c}^{(1)}-\begin{pmatrix}x_{2}^{(1)}-c_{4}^{(1)}\cdot x_{3}^{(1)}\\ -x_{1}^{(1)}+x_{3}^{(1)}\\ x_{1}^{(1)}\end{pmatrix}.$
Producing a polynomial with leading monomial $y_{2}$ is straight-forward, we simply have to take $g_{3}^{(1)}-\aleph\cdot g_{2}^{(1)}$. On the other hand, there is second polynomial with leading monomial $x_{2}^{(1)}$ after inversion of the matrices
$g_{2}^{(2)}=x_{2}^{(1)}+\left(\mathbf{A}_{2}^{-1}\mathbf{c}^{(2)}\right)_{2}-\left(-x_{1}^{(2)}+x_{3}^{(2)}\right).$
@@
$g_{1}^{(1)}+g_{2}^{(2)}=\aleph+\left(\mathbf{A}_{1}^{-1}\mathbf{c}^{(1)}\right)_{1}+c_{4}^{(1)}\cdot x_{3}^{(1)}+\left(\mathbf{A}_{2}^{-1}\mathbf{c}^{(2)}\right)_{2}+x_{1}^{(2)}-x_{3}^{(2)},$
and by the choice of our DRL term order and $c_{4}^{(1)}\neq 0$ this polynomial has leading monomial $x_{3}^{(1)}$. Obviously, after Gaussian elimination of $\mathcal{L}$ the terms of $g_{1}^{(1)}+g_{2}^{(2)},g_{2}^{(1)},g_{3}^{(1)}-\aleph\cdot g_{2}^{(1)}$ after the leading monomials will also only depend on the variables $x_{3}^{(2)},\dots,x_{3}^{(r-1)},x$.
@@
$y_{1}+\mathcal{A}_{1}\left(x_{3}^{(2)},\dots,x_{3}^{(r-1)},x\right)$ $=0,$
$y_{2}+\mathcal{A}_{2}\left(x_{3}^{(2)},\dots,x_{3}^{(r-1)},x\right)$ $=0,$
$x_{3}^{(1)}+\mathcal{A}_{3}\left(x_{3}^{(2)},\dots,x_{3}^{(r-1)},x\right)$ $=0,$
$x_{1}^{(i-1)}+c_{4}^{(i)}\cdot x_{3}^{(i)}+\ldots$ $=0,$
$x_{2}^{(i-1)}-x_{3}^{(i)}+\ldots$ $=0,$
$x_{1}^{(r-1)}+c_{4}^{(r)}\cdot x+\ldots$ $=0,$
$x_{2}^{(r-1)}-x+\ldots$ $=0,$
where $\mathcal{A}_{1}$, $\mathcal{A}_{2}$ and $\mathcal{A}_{3}$ are affine polynomials and $12\leq i\leq r-$ 1. So, for $2\leq i\leq r$ we can directly conclude from Equations 28 and 29 that
$\hat{g}_{3}^{(i)}=\left(g_{3}^{(i)}\mod(\mathcal{L})\right)\in\mathbb{F}_{q}\Big{[}x_{3}^{(2)},\dots,x_{3}^{(r-1)},x\Big{]},$
has leading monomial ${x_{3}^{(i)}}^{2}$ and respectively $x^{2}$ for the last round.
@@
Note that if we only keep the $r-1$ quadratic polynomials of $\mathcal{G}$, then we still have a fully determined polynomial system in $r-1$ variables. Therefore, we call the quadratic polynomials of $\mathcal{G}$ the downsized Ciminion polynomial system.
@@
$f_{i}$ $=x_{i}-\tilde{f}_{n}(x_{n}),\qquad 1\leq i\leq n-1,$ (30)
$f_{n}$ $=\tilde{f}_{n}(x_{n})$ (31)
It is well-known that the LEX Gröbner basis of a radical ideal can be brought into shape position for some ordering of the variables, see [KR00, Theore 3.7.25].
@@
In the situation of Theorem 3.1, let $\mathcal{G}_{\texttt{Ciminion}}$ denote the Ciminion DRL Gröbner basis.
@@
    $\dim_{\mathbb{F}_{q}}\left(\mathcal{F}_{\texttt{Ciminion}}\right)=2^{r-1}=2^{r_{C}+r_{E}-1}$.
@@
    Let $I\subset\overline{\mathbb{F}_{q}}\Big{[}y_{1},y_{2},\mathbf{x}^{(1)},\dots,\mathbf{x}^{(r-1)},x\Big{]}[x_{0}]$ be a homogeneous ideal such that $\mathcal{Z}_{+}(I)\neq\emptyset$ and $\mathcal{G}_{\texttt{Ciminion}}^{\text{hom}}\subset I$. Then $I$ is in generic coordinates.
@@
    $\operatorname{sd}_{DRL}\left(\mathcal{G}_{\texttt{Ciminion}}\right)\leq r=r_{C}+r_{E}$.
@@
    The LEX Gröbner basis for Ciminion under $y_{1}>y_{2}>\mathbf{x}^{(1)}>\ldots>\mathbf{x}^{(r-1)}>x$ is in $x$-shape position.
@@
(1), (2) and (3) are consequences of the DRL Gröbner basis from Theorem 3.1 and application of [Ste24, Theorem 3.2] and Theorem 2.6. For the last claim we take a look at the shape of the quadratic polynomials in the DRL Gröbner basis
$g_{3}^{(2)}$ $=\mathcal{Q}_{2}\left(x_{3}^{(2)},\dots,x_{3}^{(r-1)},x\right)+\mathcal{A}_{2}\left(x_{3}^{(2)},\dots,x_{3}^{(r-1)},x\right),$
$g_{3}^{(i)}$ $=\mathcal{Q}_{i}\left(x_{3}^{(i)},\dots,x_{3}^{(r-1)},x\right)+\mathcal{A}_{i}\left(x_{3}^{(i-1)},\dots,x_{3}^{(r-1)},x\right),\qquad 3\leq i\leq r$
$g_{3}^{(r)}$ $=\mathcal{Q}\left(x\right)+\mathcal{A}_{r}\left(x_{3}^{(r-1)},x\right),$
where the $\mathcal{Q}_{i}$’s are quadratic polynomials which contain the monomial ${x_{3}^{(i)}}^{2}$ and respectively $x^{2}$ for the last round, and the $\mathcal{A}_{i}$’s are affine polynomials that contain the monomial $x_{3}^{(i-1)}$ for all $i\geq 3$ and without any restrictions on $\mathcal{A}_{2}$. Then, by iterative substitution of $g_{3}^{(r)}$ into $g_{3}^{(r-1)}$ to eliminate $x_{3}^{(r-1)}$, $g_{3}^{(r-1)}$ into $g_{3}^{(r-2)}$ to eliminate $x_{3}^{(r-2)}$, etc.  we produce a LEX Gröbner basis in $x$-shape position. ∎